# Tracing Galaxy Evolution in the Nearby Universe: The Role of Dark Matter

### v1

@@ Abstract
Using a sample of $\sim$126,000 late-type galaxies (LTGs) from SDSS, we analyzed stellar mass as a function of dynamical mass. Stellar masses were estimated using eight SPS models with constant IMFs, while dynamical masses were derived from seven formulations based on Newtonian dynamics and virial equilibrium, incorporating both stellar and gas velocity dispersions. We account for key factors affecting dynamical mass estimation, including inclination, colour, concentration, and Sérsic index. We find that the difference between dynamical and stellar mass ($\Delta\log\mathbf{M}$) ranges from nearly zero to $\sim$95% of the dynamical mass, depending on mass and redshift. $\Delta\log\mathbf{M}$ appears to decreases with increasing redshift, but exhibits a saddle-like shape at low mass and low redshift—especially in disk-dominated LTGs—transitioning into a steep, linear trend at higher masses and redshifts. In the high-mass regime, the behavior resembles that of early-type galaxies. Moreover, our results indicate that this evolution is not discrete but follows a continuous transition between morphological regimes. Dark matter within LTGs is at most equal to $\Delta\log\mathbf{M}$, depending on the impact of the IMF and SPS on stellar mass estimation. Although SPS-based stellar masses do not include the gas component, previous studies have shown that galaxies with log($\mathbf{M_{Stellar}/M_{Solar}})>10$ at $z\leq 0.3$ are predominantly stellar-mass dominated. Most galaxies in our sample fall within this regime, minimizing the impact of gas exclusion. Our findings go beyond the scope of single galaxies, providing insight into the nearby Universe and highlighting the influence of dark matter in determining the Universe’s structure and evolution.
@@ I Introduction
  • Dynamical Masses: Dynamical masses are calculated using dynamic techniques that rely on rotation curves or the velocity dispersion of gas or stars within galaxies (Nigoche-Netro et al., 2022). It is important to note that these methods assume the validity of Newtonian gravity at galactic scales. Determinations of dynamical masses for late-type galaxies (LTGs) are primarily derived from their rotation curves. Early attempts to determine galaxy masses using this method were made by Scheiner (1899), Slipher (1914), Pease (1918), and Opik (1922), where they inferred the mass of M31. Later, there were efforts to determine the mass of the Milky Way by Kapteyn and van Rhijn (1922), Oort (1932a), and Oort (1932b). Determinations of M31’s total mass using velocities derived from absorption lines were conducted by Babcock (1939), Mayall (1951), and Lallemand et al. (1960). The flatness of rotation curves in various types of galaxies is a well-established phenomenon, as demonstrated by Faber and Gallagher (1979); Rubin et al. (1985); Sofue and Rubin (2001). Mass determinations of LTGs have also been made using emission lines such as $H_{\alpha}$, CO, and HI, with reasonable agreement among them.
  • Stellar, Luminous, and Baryonic Masses: In this work, we distinguish between stellar mass, luminous mass, and baryonic mass. The stellar mass refers to the mass contained in stars, whereas the luminous mass denotes the portion of the baryonic content directly traced by electromagnetic emission, predominantly from stars but potentially also from luminous gas components. The baryonic mass encompasses all baryonic components, including both luminous matter and non-luminous or weakly luminous constituents such as cold atomic and molecular gas, ionized gas, and stellar remnants.
    The luminous mass of a galaxy can be inferred from its integrated luminosity in combination with an appropriate mass-to-light ratio (M/L). The M/L ratio is not universal; it varies among galaxies and depends on factors such as mass assembly history (De Lucia et al., 2007), spectral energy distribution (SED) (Walcher et al., 2011; Conroy, 2013), stellar population mix (Oort, 1926; Baade, 1944; Van Dokkum, 2008; Conroy et al., 2009; Maraston et al., 2009; Greggio and Renzini, 2011; Maraston and Strömbäck, 2011; Chen et al., 2012), star formation rate (SFR) (Sandage, 1986; MacArthur et al., 2004), and initial mass function (IMF) (Salpeter, 1955; Scalo, 1986; Kroupa, 2001; Chabrier, 2003; Van Dokkum, 2008; Cappellari et al., 2012; Martín-Navarro et al., 2023). Although determining the M/L ratio is a non-trivial task, once it is constrained, the luminous mass can be computed directly from the galaxy’s total luminosity and its distance.
@@
    It is important to note that both the luminous and stellar masses generally underestimate the total baryonic mass, as they neglect the contribution from non-stellar baryons stored in the interstellar medium and in other non-luminous or weakly luminous components (see Section V.4).
  • Mass Determination Using Relativistic Gravitational Lensing: This method provides a rough estimate of a single galaxy’s mass. However, it usually requires a massive object like a galaxy cluster to gravitationally lens light. The first detection of weak gravitational lensing was achieved by Tyson et al. (1984) and Brainerd et al. (1996). Strong Gravitational Lensing has also been observed, with its byproducts including measurements of galaxy cluster masses Meneghetti et al. (2017), and Ghosh et al. (2023) among others.
@@
In the case of LTGs, these contain different structures (bulge, disk, spiral arms) with different levels of importance that make the proper calculation of dynamical mass difficult because the tracers of these structures are different. The best way to obtain the dynamical mass of LTGs is by using rotation curves. At low redshift, surveys such as SAMI (Sydney-AAO Multi-object Integral Field Spectrograph) and MaNGA (Mapping Nearby Galaxies at Apache Point Observatory) have provided valuable datasets. SAMI covers a redshift range of $0.004<z<0.095$ and includes approximately 3,400 galaxies across diverse environments (Croom et al., 2012), while MaNGA, part of the Sloan Digital Sky Survey (SDSS-IV), spans a redshift range of $0.01<z<0.15$ and contains a sample of around 10,000 galaxies (Bundy et al., 2015). At higher redshift, integral field units (IFUs) such as MUSE, KMOS, and SINFONI have made it possible to derive rotation curves up to $z\sim 3$, providing samples of order $\sim 1,000$ galaxies (Förster Schreiber et al., 2006; Bacon et al., 2015; Wisnioski et al., 2015). These surveys have opened a new window into the kinematics of distant LTGs. However, despite their scientific importance, the relatively small number of galaxies of these samples limit their use for statistical studies such as the one presented here, where large and homogeneous samples are required to minimize biases and obtain robust statistical inferences. In contrast, the Sloan Digital Sky Survey (SDSS) main sample contains more than one hundred thousand late-type galaxies (LTGs) within $0.00<z<0.35$, with both photometric and spectroscopic information. Although rotation curves are not available for this large sample, their dynamical masses can be estimated through the velocity dispersion of gas and/or stars. Nigoche-Netro et al. (2022) exploited this information by calibrating dynamical masses derived from SDSS velocity dispersions with rotation curves from a smaller set of LTGs in the literature, thereby accounting for the different structural components (bulge, disk, arms) that characterize LTGs.
An important feature of the rotation curves of many spiral galaxies is that in their outer part (beyond where the bulk of the galaxy’s luminosity is located), they are flat. This led to the idea, using Newton’s mechanics, that there was much more matter than what could be photometrically detected (Babcock, 1939). This non-observable matter was called dark matter. Currently, there is still no direct evidence of dark matter, and many particles have been proposed theoretically to justify it, including theories in which dark matter is dispensed with by modifying Newton’s dynamics (MOND) (Milgrom, 1983, 1988). However, these theories, like MOND, can only explain very specific phenomena, so they cannot yet be considered valid (Chan and Law, 2023, e.g.).
@@
In this work, we adopt the stellar mass as a proxy for the baryonic mass. This assumption simplifies the analysis but carries important implications: it underestimates the true baryonic mass because it neglects non-stellar baryonic components such as cold atomic and molecular gas, ionized gas, and stellar remnants. As a result, the inferred dark matter content may be correspondingly overestimated, particularly in gas-rich galaxies (details are provided in Section V.4).
@@
This logarithmic mass difference offers several methodological advantages over the commonly used dark matter fraction. It compresses the dynamical range of the measurements, improving the visibility of systematic trends and reducing the influence of outliers. As a scale-independent and symmetric estimator, it enables consistent comparisons across galaxies spanning a wide range of masses, redshifts, and environments, and integrates naturally into log-log frameworks frequently employed in galaxy scaling relations. Furthermore, it is less sensitive to absolute uncertainties in mass estimates, particularly in regimes where $\mathbf{M_{\mathbf{Dyn}}}\sim\mathbf{M_{\mathbf{Stellar}}}$, and has served as a standard diagnostic in our previous work, ensuring methodological continuity.
@@
allowing straightforward transformations between both quantities. This formulation enables direct comparison with previous results in the literature that employ either metric, while preserving the analytical and interpretive advantages of the logarithmic approach. It should be noted, however, that the general form of Equation 3 must include the total baryonic mass rather than the stellar mass alone. Consequently, when stellar mass is used as a proxy, Equation 3 is strictly valid only for systems in which non-stellar baryonic components (such as the gas mass) are negligible.
@@ II The sample of LTGs
In Figures 1, 2, and 3, we present the colour grg-r, the concentration index (R90/R50) and the Sérsic index ($n$), respectively, for the selected sample.
@@
The grg-r colour was derived from the corrected $g$ and $r$ magnitudes in the SDSS (see Section A.11).
@@ III The Mass Estimations of LTGs
In this work dynamical masses are those calculated from the velocity dispersion of gas or stars in galaxies (refer to section III.1); these were calibrated using masses derived from rotation curves. It is crucial to emphasize that velocity dispersion and rotation curves are two different methods used for mass determination. Both mass values originate from the galaxy’s gravitational potential and are calculated using Newtonian dynamics. Obtaining the dynamical mass of late-type galaxies (LTGs) is most effectively achieved through rotation curves. While integral field spectroscopy has enabled such measurements up to $z\sim 3$, the limited sample sizes and observational demands of IFU-based surveys constrain their applicability for large-scale statistical analyses. In contrast, photometric and spectroscopic data from wide-field surveys such as SDSS offer extensive coverage—over 100,000 LTGs—though they lack the spatial resolution required to derive full rotation curves. Nevertheless, dynamical masses can be estimated from gas and/or stellar velocity dispersions, providing a practical alternative for constructing homogeneous and statistically robust samples. This study leverages SDSS data to generate dynamical mass estimates for LTGs with broad coverage and relatively low associated uncertainties. By relying on spectroscopic measurements such as gas and stellar velocity dispersions, this approach enables the construction of statistically robust samples while minimizing systematic biases inherent to more spatially resolved methods. In particular, it avoids region-selection biases that arise when dynamical estimates are restricted to specific galactic zones—such as central bulges or star-forming disks—potentially misrepresenting the global mass distribution. Additionally, the use of homogeneous SDSS spectroscopic data helps mitigate instrumental and calibration biases, ensuring consistency across measurements and reducing uncertainties linked to aperture effects and cross-instrument variability.
@@
In the case of stellar mass it is derived using SPS techniques (see section III.3). It is important to recognize that SPS models typically misjudge the quantities of neutral and ionized gas, dust and stellar remnants, and other possible forms of baryonic matter in galaxies such as compact halo objects. These factors can result in an underestimation of a galaxy’s baryonic mass. Additionally, the ratios of stars, gas, and stellar remnants can differ across galaxies, influencing their dynamic characteristics and evolution. In Section V, we will examine the possible implications of using this stellar mass as an estimate for total baryonic mass in our findings.
@@
Below, we present the key variables, potential biases, and corrections that may affect the estimation of dynamical, stellar, and dark masses. A detailed description of these biases, along with the specific corrections applied to mitigate them, is provided in Appendix A, where each factor is discussed in depth and the corresponding correction formulae are outlined following Nigoche-Netro et al. (2015) and related works.
@@
  • Key variables: Velocity dispersion $\sigma$, effective radius $r_{e}$, initial mass functions IMF, colour indices, redshift $z$, stellar population models, virial factor $K$ (See equation 4), luminosity/magnitude, mass.
  • Main biases:
@@
III.1 The dynamical mass from the stars or gas of LTGs.
For the dynamical mass we adopt the principles of Newtonian dynamics, virial equilibrium and attribute velocity dispersion to the size of the gravitational potential well. Our mass estimates for LTGs are derived from data pertaining to stars and gas, sourced from the SDSS DR16. These estimates have been fine-tuned through a dataset of galaxies for which we have conducted rotation curve measurements, as detailed in Nigoche-Netro et al. (2022) (see section III.2 for more details). In SDSS DR16, velocity dispersion measurements for both stars and gas are derived from single-fiber spectroscopy obtained with the original SDSS spectrograph (used in SDSS-I/II) and the BOSS spectrograph (used in SDSS-III/IV). These instruments provide moderate-resolution spectra (R $\sim$ 1800–2200) over the wavelength range 3600–10400 Å. For stellar velocity dispersion, the SDSS DR16 pipeline employs full-spectrum fitting techniques on the observed galaxy spectra. The pipeline compares each spectrum to a grid of stellar population synthesis models (Bruzual and Charlot, 2003; Maraston, 2005), which are broadened by a Gaussian line-of-sight velocity distribution (LOSVD). The best-fit model yields an estimate of the stellar velocity dispersion, after correcting for instrumental resolution. These measurements are performed using a chi-squared minimization over selected wavelength regions that avoid strong emission lines and sky residuals. The gas velocity dispersion is calculated by fitting Gaussian profiles to prominent nebular emission lines (Balmer lines). For each line, the width of the fitted Gaussian is corrected for instrumental broadening using the known line-spread function (LSF) of the spectrograph at the corresponding observed wavelength. Emission-line measurements are performed using the idlspec2d and SpecLine pipelines, which fit both the continuum and emission-line components simultaneously or iteratively. These measurements are made available in the DR16 SpecObjAll and SpecLineAll catalogs, which include velocity dispersions, line widths, and associated uncertainties for over 4 million unique spectra.

### v2

@@ Abstract
Using a sample of $\sim$126,000 late-type galaxies (LTGs) from the Sloan Digital Sky Survey, we analyzed stellar mass as a function of the dynamical mass. Stellar masses are estimated using eight stellar population synthesis (SPS) models with constant initial mass functions (IMFs), while dynamical masses are derived from seven formulations based on Newtonian dynamics and virial equilibrium, incorporating both stellar and gas velocity dispersions. We account for key factors affecting dynamical mass estimation, including the inclination, color, concentration, and Sérsic index.
We find that the difference between dynamical and stellar mass ($\Delta\log\mathbf{M}$) ranges from nearly zero to $\sim$95% of the dynamical mass, depending on mass and redshift. $\Delta\log\mathbf{M}$ appears to decrease with increasing redshift, but exhibits a saddlelike shape at low mass and low redshift—especially in disk-dominated LTGs—transitioning into a steep, linear trend at higher masses and redshifts. In the high-mass regime, the behavior resembles that of early-type galaxies.
The results suggest that late type galaxies of higher mass or located at relatively higher redshifts appear to be more baryon-dominated, with a greater proportion of baryonic matter compared to dark matter. This can be attributed to factors such as efficient baryon condensation in the compact, gas-rich galaxies of the relatively early Universe, less concentrated dark matter halos at higher redshifts, and dynamic gas flows that accumulate baryons at galaxy centers, reducing the influence of dark matter on their dynamics.
Moreover, our results indicate that this evolution is not discrete but follows a continuous transition between morphological regimes. Dark matter within LTGs is at most equal to $\Delta\log\mathbf{M}$, depending on the impact of the IMF and SPS on stellar mass estimation. Although SPS-based stellar masses do not include the gas component, previous studies have shown that galaxies with log($\mathbf{M_{Stellar}/M_{Solar}})>10$ at $z\leq 0.3$ are predominantly stellar-mass dominated. Most galaxies in our sample fall within this regime, minimizing the impact of gas exclusion. Our findings go beyond the scope of individual galaxies, providing insight into the nearby Universe and highlighting the role of dark matter in determining galaxies’ structure and evolution.
@@ I Introduction
  • Dynamical Masses: Dynamical masses are calculated using dynamic techniques that rely on rotation curves or the velocity dispersion of gas or stars within galaxies (Nigoche-Netro et al., 2022). It is important to note that these methods assume the validity of Newtonian gravity at galactic scales. Determinations of dynamical masses for late-type galaxies (LTGs) are primarily derived from their rotation curves. Early attempts to determine galaxy masses using this method were made by Scheiner (1899), Slipher (1914), Pease (1918), and Opik (1922), where they inferred the mass of M31. Later, there were efforts to determine the mass of the Milky Way by Kapteyn and van Rhijn (1922), Oort (1932a), and Oort (1932b). Determinations of M31’s total mass using velocities derived from absorption lines were conducted by Babcock (1939), Mayall (1951), and Lallemand et al. (1960). The flatness of rotation curves in various types of galaxies is a well-established phenomenon, as demonstrated by Faber and Gallagher (1979); Rubin et al. (1985); Sofue and Rubin (2001). Mass determinations of LTGs have also been made using emission lines such as Hα, CO, and HI, with reasonable agreement among them.
  • Stellar, Luminous, and Baryonic Masses: In this work, we distinguish between stellar mass, luminous mass, and baryonic mass. The stellar mass refers to the mass contained in stars, whereas the luminous mass denotes the portion of the baryonic content directly traced by electromagnetic emission, predominantly from stars but potentially also from luminous gas components. The baryonic mass encompasses all baryonic components, including both luminous matter and nonluminous or weakly luminous constituents such as cold atomic and molecular gas, ionized gas, and stellar remnants.
    The luminous mass of a galaxy can be inferred from its integrated luminosity in combination with an appropriate mass-to-light ratio (M/L). The M/L ratio is not universal; it varies among galaxies and depends on factors such as mass assembly history (De Lucia et al., 2007), spectral energy distribution (SED, Walcher et al., 2011; Conroy, 2013), stellar population mix (Oort, 1926; Baade, 1944; Van Dokkum, 2008; Conroy et al., 2009; Maraston et al., 2009; Greggio and Renzini, 2011; Maraston and Strömbäck, 2011; Chen et al., 2012), star formation rate (SFR, Sandage, 1986; MacArthur et al., 2004), and initial mass function (IMF, Salpeter, 1955; Scalo, 1986; Kroupa, 2001; Chabrier, 2003; Van Dokkum, 2008; Cappellari et al., 2012; Martín-Navarro et al., 2023). Although determining the M/L ratio is a nontrivial task, once it is constrained, the luminous mass can be computed directly from the galaxy’s total luminosity and its distance.
@@
    It is important to note that both the luminous and stellar masses generally underestimate the total baryonic mass, as they neglect the contribution from nonstellar baryons stored in the interstellar medium and in other nonluminous or weakly luminous components (see Section V.4).
  • Mass Determination Using Relativistic Gravitational Lensing: This method provides a rough estimate of a single galaxy’s mass. However, it usually requires a massive object like a galaxy cluster to gravitationally lens light. The first detection of weak gravitational lensing was achieved by Tyson et al. (1984) and Brainerd et al. (1996). Strong gravitational lensing has also been observed, with its byproducts including measurements of galaxy cluster masses, by Meneghetti et al. (2017), and Ghosh et al. (2023) among others.
@@
In the case of LTGs, these contain different structures (bulge, disk, spiral arms) with different levels of importance that make the proper calculation of dynamical mass difficult because the tracers of these structures are different. The best way to obtain the dynamical mass of LTGs is by using rotation curves. At low redshift, surveys such as SAMI (Sydney-AAO Multi-object Integral Field Spectrograph) and MaNGA (Mapping Nearby Galaxies at Apache Point Observatory) have provided valuable datasets. SAMI covers a redshift range of $0.004<z<0.095$ and includes approximately 3400 galaxies across diverse environments (Croom et al., 2012), while MaNGA, part of the Sloan Digital Sky Survey (SDSS-IV), spans a redshift range of $0.01<z<0.15$ and contains a sample of around 10,000 galaxies (Bundy et al., 2015). At higher redshift, integral field units (IFUs) such as MUSE, KMOS, and SINFONI have made it possible to derive rotation curves up to $z\sim 3$, providing samples of order $\sim 1000$ galaxies (Förster Schreiber et al., 2006; Bacon et al., 2015; Wisnioski et al., 2015). These surveys have opened a new window into the kinematics of distant LTGs. However, despite their scientific importance, the relatively small number of galaxies in these samples limits their use for statistical studies such as the one presented here, where large and homogeneous samples are required to minimize biases and obtain robust statistical inferences. In contrast, the Sloan Digital Sky Survey (SDSS) main sample contains more than one hundred thousand late-type galaxies (LTGs) within $0.00<z<0.35$, with both photometric and spectroscopic information. Although rotation curves are not available for this large sample, their dynamical masses can be estimated through the velocity dispersion of gas and/or stars. Nigoche-Netro et al. (2022) exploited this information by calibrating dynamical masses derived from SDSS velocity dispersions with rotation curves from a smaller set of LTGs in the literature, thereby accounting for the different structural components (bulge, disk, arms) that characterize LTGs.
An important feature of the rotation curves of many spiral galaxies is that in their outer part (beyond where the bulk of the galaxy’s luminosity is located), they are flat. This led to the idea, using Newton’s mechanics, that there was much more matter than what could be photometrically detected (Babcock, 1939). This nonobservable matter was called dark matter. Currently, there is still no direct evidence of dark matter, and many particles have been proposed theoretically to justify it, including theories in which dark matter is dispensed with by modifying Newton’s dynamics (MOND, Milgrom, 1983, 1988). However, these theories, like MOND, can only explain very specific phenomena, so they cannot yet be considered valid (e.g. Chan and Law, 2023).
@@
In this work, we adopt the stellar mass as a proxy for the baryonic mass. This assumption simplifies the analysis but carries important implications: it underestimates the true baryonic mass because it neglects nonstellar baryonic components such as cold atomic and molecular gas, ionized gas, and stellar remnants. As a result, the inferred dark matter content may be correspondingly overestimated, particularly in gas-rich galaxies (details are provided in Section V.4).
@@
This logarithmic mass difference offers several methodological advantages over the commonly used dark matter fraction. It compresses the dynamical range of the measurements, improving the visibility of systematic trends and reducing the influence of outliers. As a scale-independent and symmetric estimator, it enables consistent comparisons across galaxies spanning a wide range of masses, redshifts, and environments, and integrates naturally into log–log frameworks frequently employed in galaxy scaling relations. Furthermore, it is less sensitive to absolute uncertainties in mass estimates, particularly in regimes where $\mathbf{M_{\mathbf{Dyn}}}\sim\mathbf{M_{\mathbf{Stellar}}}$, and has served as a standard diagnostic in our previous work, ensuring methodological continuity.
@@
allowing for straightforward transformations between both quantities. This formulation enables direct comparison with previous results in the literature that employ either metric, while preserving the analytical and interpretive advantages of the logarithmic approach. It should be noted, however, that the general form of Equation 3 must include the total baryonic mass rather than the stellar mass alone. Consequently, when stellar mass is used as a proxy, Equation 3 is strictly valid only for systems in which nonstellar baryonic components (such as the gas mass) are negligible.
@@ II The sample of LTGs
In Figures 1, 2, and 3, we present the color grg-r, the concentration index (R90/R50) and the Sérsic index ($n$), respectively, for the selected sample.
@@
The grg-r color was derived from the corrected $g$ and $r$ magnitudes in the SDSS (see Appendix A.11).
@@ III The Mass Estimations of LTGs
In this work, dynamical masses are those calculated from the velocity dispersion of gas or stars in galaxies (refer to section III.1); these were calibrated using masses derived from rotation curves. It is crucial to emphasize that velocity dispersion and rotation curves are two different methods used for mass determination. Both mass values originate from the galaxy’s gravitational potential and are calculated using Newtonian dynamics. Obtaining the dynamical mass of late-type galaxies (LTGs) is most effectively achieved through rotation curves. While integral field spectroscopy has enabled such measurements up to $z\sim 3$, the limited sample sizes and observational demands of IFU-based surveys constrain their applicability for large-scale statistical analyses. In contrast, photometric and spectroscopic data from wide-field surveys such as SDSS offer extensive coverage—over 100,000 LTGs—though they lack the spatial resolution required to derive full rotation curves. Nevertheless, dynamical masses can be estimated from gas and/or stellar velocity dispersions, providing a practical alternative for constructing homogeneous and statistically robust samples. This study leverages SDSS data to generate dynamical mass estimates for LTGs with broad coverage and relatively low associated uncertainties. By relying on spectroscopic measurements such as gas and stellar velocity dispersions, this approach enables the construction of statistically robust samples while minimizing systematic biases inherent to more spatially resolved methods. In particular, it avoids region-selection biases that arise when dynamical estimates are restricted to specific galactic zones—such as central bulges or star-forming disks—potentially misrepresenting the global mass distribution. Additionally, the use of homogeneous SDSS spectroscopic data helps mitigate instrumental and calibration biases, ensuring consistency across measurements and reducing uncertainties linked to aperture effects and cross-instrument variability.
@@
In the case of stellar mass, it is derived using SPS techniques (see section III.3). It is important to recognize that SPS models typically misjudge the quantities of neutral and ionized gas, dust, and stellar remnants, and other possible forms of baryonic matter in galaxies such as compact halo objects. These factors can result in an underestimation of a galaxy’s baryonic mass. Additionally, the ratios of stars, gas, and stellar remnants can differ across galaxies, influencing their dynamic characteristics and evolution. In Section V, we will examine the possible implications of using this stellar mass as an estimate for total baryonic mass in our findings.
@@
Below, we present the key variables, potential biases, and corrections that may affect the estimation of dynamical, stellar, and dark masses. A detailed description of these biases, along with the specific corrections applied to mitigate them, is provided in Appendix A, where each factor is discussed in depth, and the corresponding correction formulae are outlined following Nigoche-Netro et al. (2015) and related works.
@@
  • Key variables. Velocity dispersion $\sigma$, effective radius $r_{e}$, initial mass functions IMF, color indices, redshift $z$, stellar population models, virial factor $K$ (see equation 4), luminosity/magnitude, mass.
  • Main biases.
@@
III.1 The dynamical mass from the stars or gas of LTGs
For the dynamical mass, we adopt the principles of Newtonian dynamics, virial equilibrium, and attribute velocity dispersion to the size of the gravitational potential well. Our mass estimates for LTGs are derived from data pertaining to stars and gas, sourced from the SDSS DR16. These estimates have been fine-tuned through a dataset of galaxies for which we have conducted rotation curve measurements, as detailed in Nigoche-Netro et al. (2022) (see section III.2 for more details). In SDSS DR16, velocity dispersion measurements for both stars and gas are derived from single-fiber spectroscopy obtained with the original SDSS spectrograph (used in SDSS-I/II) and the BOSS spectrograph (used in SDSS-III/IV). These instruments provide moderate-resolution spectra (R $\sim$ 1800–-2200) over the wavelength range 3600–-10400 Å. For stellar velocity dispersion, the SDSS DR16 pipeline employs full-spectrum fitting techniques on the observed galaxy spectra. The pipeline compares each spectrum to a grid of stellar population synthesis models (Bruzual and Charlot, 2003; Maraston, 2005), which are broadened by a Gaussian line-of-sight velocity distribution (LOSVD). The best-fit model yields an estimate of the stellar velocity dispersion, after correcting for instrumental resolution. These measurements are performed using a chi-squared minimization over selected wavelength regions that avoid strong emission lines and sky residuals. The gas velocity dispersion is calculated by fitting Gaussian profiles to prominent nebular emission lines (Balmer lines). For each line, the width of the fitted Gaussian is corrected for instrumental broadening using the known line-spread function (LSF) of the spectrograph at the corresponding observed wavelength. Emission-line measurements are performed using the idlspec2d and SpecLine pipelines, which fit both the continuum and emission-line components simultaneously or iteratively. These measurements are made available in the DR16 SpecObjAll and SpecLineAll catalogs, which include velocity dispersions, line widths, and associated uncertainties for over 4 million unique spectra.